-3(x+4)-5(x-4)


9/16-3/8+1/4

2(x-7)-5=5-(3-2x)

Answer by rfer(16322)

-3x-12-5x+20=
-8x+8
---------------
9/16-6/16+4/16=7/16
------------------
2x-14-5=5-3+2x=21